# Scaling Law with Learning Rate Annealing

### v1

@@ Abstract
We find that the cross-entropy loss curves of neural language models empirically adhere to a scaling law with learning rate (LR) annealing over training steps ($s$):
$L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$
Where $S_{1}$ is forward area and $S_{2}$ is learning rate annealing area. This formulation takes into account two factors: (1) The forward scaling defined as typical scaling law, and (2) the additional loss drop brought by LR annealing. Therefore, this formulation can describe the full loss curve at each step, rather than the single loss point at the end of training. Applying the scaling law with LR annealing and fitting only one or two training curves, we can accurately predict the loss of language model training at any given step and across any learning rate scheduler (LRS). Furthermore, this equation accurately describes the dynamics during training process, and provides a theoretical verification and explanation for numerous experimental findings of previous studies, particularly those focusing on LR schedule and LR annealing. The resulting insights, also serve as a guide for researchers to select critical LRS in advance by prediction using our equation. Most significantly, since all the points in a full training curve follow the equation, we can achieve accurate loss prediction at any given step across any learning rate scheduler, while expending less than 1% of the computational cost required by the chinchilla scaling law to fit language modeling loss. This approach extremely democratizes scaling law fitting and predicting in developing large language models.
@@ 1 Introduction
In recent years, the spotlight of academic and industrial interest has been increasingly captured by large language models. The intriguing notion of the scaling law posits that the cross-entropy loss of language models adheres to a power-law pattern as scaling up of model size and data size (Kaplan et al., 2020; Hoffmann et al., 2022). This law serves as a powerful tool, affording researchers the ability to precisely forecast the performance of large language models by fitting the loss at a smaller scale. The potential implications of scaling laws are vast and compelling, sparking a surge of exploration in this burgeoning field (Bahri et al., 2021; Michaud et al., 2023; DeepSeek-AI, 2024).
However, typical scaling laws describe the performance of the model at the end of training, rather than every step during the training process. Essentially, the middle points with different degrees of LR annealing fail to follow typical scaling laws, which do not consider local loss drop brought by LR annealing. Consequently, these laws are unable to fit or predict a full loss curve. Till this work, we do not have an appropriate formulation that accurately describes the dynamics during the training process, which is crucial to deeply understand and improve the training process.
The learning rate schedule (LRS) is a significant variable that greatly influences the loss during the entire training process. Several previous studies have concluded that the rate of decrease in loss accelerates as LR annealing takes place (Loshchilov & Hutter, 2016; Ibrahim et al., 2024; Hu et al., 2024; DeepSeek-AI, 2024). However, these studies generally only provide a qualitative description of how loss changes during LR annealing. Instead, in the present study, we build upon the chinchilla scaling law (Hoffmann et al., 2022) by introducing an additional LR annealing term. This term allows us to accurately predict how loss decreases during LR annealing. Overall, we discover that the loss value at any step of a fixed-size model is determined by two factors: forward area and the degree of LR annealing. Formally, the expectation of loss $L$ at step  [^1] In this paper, we take training steps as data amount in the scaling law equation, which has no essential difference because data amount $=$ training steps $\times$ batch size and batch size is fixed. $s$ of a language model follows:
$L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$ (1)
$S_{1}=\sum\limits_{i=1}^{s}\eta_{i}$
$S_{2}=\sum\limits_{i=1}^{s}\sum\limits_{k=1}^{i}(\eta_{k-1}-\eta_{k})\cdot\lambda^{i-k}$
[…]
Figure 1: The definition of $S_{1}$ and $S_{2}$, taking a 20-step cosine learning rate schedule as a toy example. $S_{1}$ (forward area) is the area enclosed by the learning rate curve and the X-axis step, which can be approximately regarded as the total amount of movement of neural network parameters; $S_{2}$ is the decayed annealing area of the learning rate curve (the lighter the color, the greater the degree of decay), that is, the weighted sum of the areas of the small blue squares. Both the increase of $S_{1}$ and $S_{2}$ can help to reduce the loss, but they exist in a delicate balance or trade-off relationship. To minimize the final loss, it’s necessary to discover optimal balance between $S_{1}$ and $S_{2}$.
@@
In Eq. 1, the term $L_{0}+A\cdot S_{1}^{-\alpha}$ is a rectified scaling law, which states that the expected loss decreases as a power-law function of the number of training steps. The term $-C\cdot S_{2}$ is the newly introduced learning rate annealing term. This term represents the additional decrease in loss that occurs as the learning rate is reduced during training. This universal equation can fit the validation loss of any step across any LRS, and then help predict the validation loss of any step across any LRS. As some examples, we fit the equation on the loss curve of constant and cosine LRS in 20K total steps (shown as Fig. 2), then we are able to employ the universal scaling law with LR annealing to predict the full loss curve across various LRS in longer total steps (e.g. 60K) (shown as Fig. 3).
In Sec. 3, we demonstrate the derivation of the scaling law with LR annealing and elucidate the potential theory underpinning our equation. We also extend our equation to a formulation with model size $N$. In Sec. 4, we apply our equation to verify and explain the conclusions drawn from a multitude of previous studies. These conclusions, which previously required substantial computational resources, can now be consolidated in a cost-free manner using our equation. This not only streamlines the process but also brings us closer to understanding the crux of everything about loss drop, LR schedule, and LR annealing (i.e., the art of balance between forward area and annealing area). Moreover, these key insights serve as a guide for researchers to select critical LRS in advance by prediction using our equation. In Sec. 5, we compare our equation with typical chinchilla scaling law (Hoffmann et al., 2022). We show that chinchilla scaling law actually describes special cases (the endpoint of one full loss curve) of our equation, which means that our equation is a generalized form of chinchilla scaling law. Furthermore, we compare the computational costs required by the chinchilla scaling law and our equation, revealing that our equation offers significant savings in computational costs. This greatly democratizes the fitting and prediction of scaling laws.
(a) Learning rate.
(b) Zoomed-out loss fitting.
(c) Zoomed-in loss fitting.
Figure 2: Full loss curve fitting on cosine (20k steps to $\eta_{min}=0$) and constant LRS. The figures omit the warmup in the first 500 steps. After fitting, we get a universal loss equation $L=2.628+0.429\cdot S_{1}^{-0.550}-0.412\cdot S_{2}$.
@@
(b) Full curve prediction of cosine LRS (60K steps to $\eta_{min}=0.1\cdot\eta_{max}$).
(c) Full curve prediction of multi-step cosine LRS (80% + 10% + 10%) (DeepSeek-AI, 2024)
(d) Full curve prediction of WSD LRS (20% cosine annealing to $\eta_{min}=0$(Hu et al., 2024).
(e) Full curve prediction of even strange LRS: periodically repeating linear annealing and linear re-warmup.
Figure 3: Full loss curve prediction (60K steps) by the universal loss curve equation across various LRS. The left, the medium, and the right figures in each row are learning rate curve, zoomed-out loss prediction, zoomed-in loss prediction, respectively. The red rectangle means the zoomed-in zone. The figures omit the warmup in the first 500 steps. The universal loss curve is $L=2.628+0.429\cdot S_{1}^{-0.550}-0.412\cdot S_{2}$, fitted as in Fig. 2. The equation can accurately predict various forms of unseen LRS, and it is highly precise in predicting the trend of loss changes as LR varies. Please note that these are predictive results, which means that none of the points in this figure are involved in the fitting process. The mean prediction errors across various LRS are low to $\sim 0.2\%$.
@@ 2.1 Scaling Laws
Cross-entropy loss of language models is a good indicator for the performance of almost all downstream tasks (Caballero et al., 2022; Du et al., 2024)Kaplan et al. (2020) empirically discovers a power-law relationship between validation loss and three key factors: model size, dataset size, and the computational resources utilized for training. As an application of scaling law, Hoffmann et al. (2022) trains a compute-optimal large language model named chinchilla by balancing model size and dataset size. Moreover, chinchilla scaling law adopts a simpler and intuitive equation to describe the relationship between the final validation loss ($L$) and the number of parameters ($N$) as well as the amount of data ($D$) as follows:
$L(D,N)=L_{0}+A\cdot D^{-\alpha}+B\cdot N^{-\beta}$ (2)
where $L_{0}$, $A$, $B$, $\alpha$, $\beta$ are all undetermined positive constants. Note that the typical scaling laws fit the loss at the endpoint of training instead of full loss curve, thus a experiment with new data size requires another training launch. Previous works have conducted some preliminary studies on the impact of the learning rate on the scaling laws. For example, OpenAI and chinchilla scaling laws both find that the choice of learning rate schedule does not influence the power-law format (Kaplan et al., 2020; Hoffmann et al., 2022). Also, OpenAI’s experiments suggest that the specific choice of learning rate schedule has minimal impact on the final validation loss, provided that the total summed learning rate is adequately large and the schedule incorporates both a warmup stage and a final annealing stage, reducing the learning rate to nearly zero (Kaplan et al., 2020).
2.2 WSD Learning Rate Scheduler
$WSD(s)=\begin{cases}&\frac{s}{T_{warmup}}\eta_{max},\quad s\leq T_{warmup}\\ &\eta_{max},\quad T_{warmup}<s\leq T_{stable}\\ &f(s-T_{stable})\eta_{max},\quad T_{stable}<s\leq T_{total}\end{cases}$ (3)
Where $0\leq f(s-T_{stable})\leq 1$ is a decreasing function about step $s$, and $\eta_{max}$ is the maximal learning rate. Hägele et al. (2024) consolidates the effectiveness of WSD scheduler by many empirical experiments. Moreover, Hägele et al. (2024) also finds that using 1-sqrt annealing and a moderate annealing ratio (e.g. 20%) can further decrease the final loss.
@@ 2.3 Learning Rate Annealing
Learning rate annealing is a common technique used in training neural networks, where the learning rate is reduced over training iterations according to a pre-defined learning rate schedule. This method is often employed to improve the performance and stability of the model during training (Loshchilov & Hutter, 2016). For example, the widely-used cosine LRS could be perceived as a process where the learning rate is cosine curve-like annealing over full steps. Conversely, WSD LRS (Hu et al., 2024) maintains a stable learning rate for the majority of the steps, implementing annealing only in the final (e.g. $10\%\sim 20\%$) of the steps. During the training of LLMs, it has been widely observed that a more pronounced decrease in the learning rate often results in a more precipitous drop in the validation loss (Loshchilov & Hutter, 2016; Ibrahim et al., 2024; DeepSeek-AI, 2024; Hu et al., 2024). However, to the best of our knowledge, all previous studies end with providing a rough and qualitative description of how loss changes during LR annealing, while we provide an equation to formulate the loss changes during LR annealing in this work.
@@ 3 Theory
In this section, we elaborate in detail on the intuition, the origin, step-by-step derivation, and the experimental basis of our theory. We then prove the effectiveness of our formula through experiments.
@@ 3.1 Similarity between Learning Rate, Gradient Norm, and Loss
Figure 4: The curve shapes of learning rate (top), gradient norm (medium), and validation loss (bottom) hold a quite high similarity across various LRS (labeled as different colors).
As shown in Fig. 4, the first key observation is that the shapes of LR curve, gradient norm curve, and validation loss curve are quite similar across various LRS. This gives a clue that there might exist an implicit connection between learning rate and loss, where gradient norm could be the bridge.
@@
Constant LRS could be seen as a special LRS, in which any step could be as the endpoint of the LRS. Memorize that chinchilla scaling law (Hoffmann et al., 2022) exactly describes the endpoint of full loss curves. That is to say, the expectation of validation loss of all steps in constant LRS adhere to chinchilla scaling law.
@@
Differently, LR annealing (or re-warmup) brings drastic loss changes in a local range (e.g. Fig. 4), which makes the full loss curve no more follow typical scaling laws. Consider that the LRS without annealing, or constant LRS, follows typical scaling laws. Based on that, we can easily induce that, in a full loss curve, the loss $L$ at step $s$ can be rectified by an added $\eta$-related item, and may follow the format as
$L(s)=\color[rgb]{0,1,1}L_{0}+A\cdot s^{-\alpha}\color[rgb]{1,0,0}-f(\eta)$ (4)
where the blue part is the same as chinchilla scaling law (Hoffmann et al., 2022), while the red part, $f(\eta)$ denotes the extra loss change brought by LR annealing. Recall the curve similarity between learning rate and loss. It could lead to one naive guess: $f(\eta)=C\cdot\eta_{s}$, where $C$ is a positive constant.
@@
The form of Eq. 4 is still imperfect. Note that the gradient $\mathbf{g}$ of parameters decreases along with LR (shown in Fig. 4). Compared with the stages before annealing, the amount of parameter changes (nearly $\eta\cdot|\mathbf{g}|$ of each step) in the neural network optimization process will bring out an almost quadratic faster decline during LR annealing. These clues show that during LR annealing, the loss drop brought by power law (the blue part in Eq. 4) should also decrease. Therefore, we modify to the following format:
@@
$S_{1}$ $=\sum\limits_{i=1}^{s}\eta_{i}$
Where we define $S_{1}$ as forward area. $S_{1}$ is exactly the area enclosed by the LRS curve and the X-axis step, which could be approximately regarded as the total amount of neural network parameter changes. Fig. 1 gives a more intuitive visualization for the definition of $S_{1}$.

### v2

@@ Abstract
We find that the cross-entropy loss curves of neural language models empirically adhere to a scaling law with learning rate (LR) annealing over training steps:
$L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2},$
where $L(s)$ is the validation loss at step $s$, $S_{1}$ is the area under the LR curve, $S_{2}$ is the LR annealing area, and $L_{0}$, $A$, $C$, $\alpha$ are constant parameters. This formulation takes into account two factors: (1) power-law scaling over data size, and (2) the additional loss reduction during LR annealing. Therefore, this formulation can describe the full loss curve at each step, rather than the single loss point at the end of training. Applying the scaling law with LR annealing and fitting only one or two training curves, we can accurately predict the loss at any given step across any learning rate scheduler (LRS). This approach significantly reduces computational cost in formulating scaling laws while providing more accuracy and expressiveness for training dynamics. Extensive experiments demonstrate that our findings hold across a range of hyper-parameters and model architectures, and our equation can extend to scaling effect of model sizes. Moreover, our formulation provides accurate theoretical verification and explanation for empirical results observed in numerous previous studies, particularly those focusing on LR schedule and annealing. We believe that this work is promising to enhance the understanding of LLM training dynamics while greatly democratizing scaling laws, and it can guide researchers in refining training strategies (e.g. critical LRS) for further LLMs  [^1] We welcome any feedback, comment, and discussion at AlphaXiv or HuggingFace, where we also provide our implementation codes..
@@ 1 Introduction
In recent years, large language models (LLMs) have garnered significant academic and industrial attention (Brown et al., 2020; Touvron et al., 2023). The scaling law suggests that the validation loss of language models follow a power-law pattern as model and data sizes increase (Hestness et al., 2017; Kaplan et al., 2020; Henighan et al., 2020). This law provides a powerful framework for forecasting LLM performances before large scale training by fitting losses at smaller scales (OpenAI, 2023; DeepSeek-AI, 2024; Dubey et al., 2024). Numerous studies have explored on the formulation to model the scaling effect of LLMs under various different settings (Bahri et al., 2021; Hernandez et al., 2021; Caballero et al., 2022; Michaud et al., 2023; Muennighoff et al., 2023).
However, typical scaling law formulations focus only on the final loss at the end of training (Hoffmann et al., 2022). Specifically, previous approaches generally rely on a set of training runs and fit the scaling law curve solely on the final loss from each run. Essentially, the middle points with different degrees of LR annealing fail to follow typical scaling laws, which do not consider local loss drop brought by LR annealing. The previous approach under-utilizes the training compute and fails to capture the training dynamics within each run. Further, the application of scaling laws in LLM developments is limited since the loss curve through the whole training process is not modeled. An expressive formulation that models full loss curves enables prediction of future training dynamics and also offers insights on understanding the learning process of LLMs.
In this study, we propose a scaling law that models the full loss curve within a complete LLM training run. Specifically, we dive deeper into the training dynamics during LR annealing, and incorporate a LR annealing factor into the traditional scaling law formula to formulate the process. This design is motivated by the observed correlation between LRS and loss curves, where loss gradually decreases as we consume more training steps  [^2] In this paper, we use training steps to quantify the amount of consumed data, as they are typically proportional, with data amount calculated as training steps multiplied by batch size. and then sharply declines when the LR undergoes significant annealing (Loshchilov & Hutter, 2016; Smith et al., 2018; Ibrahim et al., 2024; Hu et al., 2024). Fig. 4 depicts how loss curves change over different learning rate schedules. Overall, we discover that the validation loss of a language model at any step is determined by two factors: the forward area $S_{1}$ under the LR curve and the degree of LR annealing $S_{2}$ at that step. Formally, the expectation of loss $L$ at step $s$ of a language model follows::
$L(s)$ $=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2},$ (1)
$S_{1}$ $=\sum\limits_{i=1}^{s}\eta_{i},$
$S_{2}$ $=\sum\limits_{i=1}^{s}\sum\limits_{k=1}^{i}(\eta_{k-1}-\eta_{k})\cdot\lambda^{i-k},$
where $\eta_{i}$ is the learning rate at step $i$, and $\lambda$ is a hyper-parameter representing the decay factor for LR annealing momentum (see Sec. 3 in detail),which typically ranges from $0.99$ to $0.999$. $L_{0}$, $A$, $C$, $\alpha$ are undetermined positive constants. $S_{1}$ is also known as the summed learning rate (Kaplan et al., 2020), and $S_{2}$ represents the LR annealing area. A visualization of $S_{1}$ and $S_{2}$ is provided as Fig. 1.
@@
Eq. 1 describes how loss changes for each step in a full loss curve during training. In Eq. 1, the term $L_{0}+A\cdot S_{1}^{-\alpha}$ represents a rectified scaling law that captures the expected loss decreases as a power-law function of the number of training steps. The new term $-C\cdot S_{2}$ accounts for the further loss drop due to learning rate annealing. Remarkably, this simple formulation accurately describes the validation loss at any training step across various LRS and even allows us to predict the loss curve for unseen LRS. For example, we can fit Eq. 1 to the full loss curve of constant and cosine LRS with 20K total steps (Fig. 2), and then predict the full loss curve for various unseen LRS with longer total steps (e.g. 60K) (Fig. 3).
Figure 1: Visualization of $S_{1}$ and $S_{2}$ at the 20-th step of a cosine LR scheduler. $S_{1}$ is the forward area, i.e., sum of red grid areas, which can be approximately regarded as the total amount of movement for neural network parameters; $S_{2}$ is the decayed annealing area, i.e., weighted sum of blue grid areas, where lighter shades indicate smaller weights. Both $S_{1}$ and $S_{2}$ contribute to loss reduction, and balancing their values is crucial for achieving the lowest possible final loss.
(a) LR curves.
(b) Loss curves.
(c) Zoomed-in view.
Figure 2: Using Eq. 1 to fit full loss curves yield by constant and cosine LRS. Total steps = 20K, $\eta_{max}=2\times 10^{-4}$, $\eta_{min}=0$. The fitted equation is $L(s)=2.628+0.429\cdot S_{1}^{-0.550}-0.411\cdot S_{2}$.
@@
(b) Full loss curve prediction of the cosine LRS (60K steps, $\eta_{min}=0.1\cdot\eta_{max}$).
(c) Full loss curve prediction of the multi-step cosine LRS (80% + 10% + 10%) (DeepSeek-AI, 2024).
(d) Full Loss curve prediction of the WSD LRS (20% cosine annealing to $\eta_{min}=0$(Hu et al., 2024).
(e) Full Loss curve prediction of the Cyclic LRS (Smith, 2017) including multiple unseen LR re-warmup stages.
Figure 3: Using the fitted equation from Fig. 2 to predict full loss curves for unseen LRS with 60K total steps. The left, middle, and right columns present the LR curve, the loss curve, and a zoomed-in view of loss curve, respectively. Warmup steps (500) are not shown in this figure. The fitted equation accurately predicts each loss curve, particularly for capturing the trend of loss changes as the LR varies. Notable, all LRS and loss curves shown here were unseen during the fitting in Fig. 2. The mean prediction errors across different LRS is as low as $\sim 0.2\%$.
We validate our proposed equation through extensive experiments and find that: (1) Our formulation performs consistently well across various hyper-parameters and model architectures; (2) Eq. 1 can be extended to incorporate other scaling factors, such as model sizes; (3) Our proposed equation accurately fits the loss curves of open-sourced models; (4) Our formulation can be used to verify and explain numerous previous findings regarding LR annealing and scheduling.
In Sec. 3, we derive the scaling law formulation with LR annealing and elucidate the potential theory underpinning our formulation. Extensive experiments are conducted to validate the formulation. In Sec. 4, we apply our formulation to verify and explain the empirical results from various previous studies. Our approach offers theoretical insights into the crux of loss drop, LR schedule, and LR annealing, enabling LLM participants to better understand training dynamics of LLM and select optimal training recipes in advance. In Sec. 5, we compare our approach to typical scaling law formula, such as the Chinchilla scaling law (Hoffmann et al., 2022). We show that our formulation is more general and requires significantly less compute (less than 1%) to fit, which greatly democratizes the development of LLMs and scaling laws.
@@ 2.1 Scaling Laws
Cross-entropy loss of language models on the validation set is a reliable indicator of LLMs’ performance on downstream tasks (Caballero et al., 2022; Du et al., 2024). Kaplan et al. (2020) empirically discovered a power-law relationship between validation loss $L$ and three factors: model size $N$, dataset size $D$, and training compute. As an application of scaling law, Hoffmann et al. (2022) developed Chinchilla, a compute-optimal LLM, by balancing model size and dataset size. They used a simplified and intuitive scaling law equation:
$L(D,N)=L_{0}+A\cdot D^{-\alpha}+B\cdot N^{-\beta},$ (2)
where $L_{0}$, $A$, $B$, $\alpha$, $\beta$ are positive constants. Traditional scaling law formulations fit only the loss at the final training step, while ignoring losses from other steps. Note that collecting a new loss value of data size requires launching a another training run with the same LRS, which is resource-intensive. Previous works have conducted some preliminary studies on the impact of the learning rate on the scaling laws. For example, OpenAI and chinchilla scaling laws both report that the choice of learning rate schedule does not influence the power-law format (Kaplan et al., 2020; Hoffmann et al., 2022). Also, OpenAI’s experiments suggest that the specific choice of learning rate schedule has minimal impact on the final validation loss, provided that the total summed learning rate is adequately large and the schedule incorporates both a warmup stage and a final annealing stage, reducing the learning rate to nearly zero at the end of training (Kaplan et al., 2020).
@@ 2.2 Learning Rate Annealing
Learning rate annealing is a widely-used technique in training neural networks, where the learning rate is progressively reduced from a maximum to a minimum value following a pre-defined LRS. Various LRS schemes have been proposed to improve the performance and stability of model training. For example, the popular cosine LRS (Loshchilov & Hutter, 2016) reduces the LR in a cosine-like pattern over full training steps. WSD LRS (Hu et al., 2024) keeps a constant LR for the majority of training, and applies annealing only in the final (e.g. $10\%\sim 20\%$) steps. During the training of LLMs, it has been widely observed that a more pronounced decrease in the learning rate often results in a more precipitous drop in the validation loss (Loshchilov & Hutter, 2016; Ibrahim et al., 2024; DeepSeek-AI, 2024; Hu et al., 2024). However, to the best of our knowledge, all previous studies end with providing a rough and qualitative description of how loss changes during LR annealing, while our work provides an accurate equation to quantitatively formulate the loss changes during LR annealing.
@@ 3 Theory
In this section, we elaborate the origin, the intuition, and the experimental basis behind Eq. 1. We then validate our formula through extensive experiments.
@@ 3.1 Similarity between Learning Rate, Gradient Norm, and Loss
Figure 4: The shapes of LR (top), gradient norm (medium), and validation loss (bottom) curves exhibit high similarity across various LRS (labeled as different colors).
The first key observation is that the shapes of LR curve, gradient norm curve, and validation loss curve are quite similar across various LRS when training LLMs (Fig. 4). This suggests an implicit connection between learning rate and loss, where gradient norm could be the bridge.
@@
A constant LRS is a special LRS, where every training step can be viewed as an endpoint of the LRS. Notably, the Chinchilla scaling law (Hoffmann et al., 2022) exactly fits losses of last steps, i.e., LRS endpoints. Therefore, the expectation of validation loss of all steps under in constant LRS adheres to a power-law over training step $s$.
@@
Unlike a constant LRS, LR annealing (or re-warmup) brings significant local changes in the loss (see Fig. 4), causing the full loss curve to deviate from the traditional power-law formulation that consider only the training steps $s$. We hypothesis that such loss changes can be captured by an additional LR ($\eta$) related term, i.e.,
$L(s)={\color[rgb]{0,1,1}\underbrace{L_{0}+A\cdot s^{-\alpha}}_{\text{Traditional scaling law}}}\quad{\color[rgb]{1,0,0}\underbrace{-f(\eta)}_{\text{LR annealing term}}},$ (3)
where the first two terms (blue part) follow traditional scaling laws, while the last term (red part) denotes the extra loss change brought by LR annealing. Recall the similarity between learning rate and loss curves, we can form a naive guess for $f(\eta)$ as $f(\eta)=C\cdot\eta_{s}$, where $C$ is a positive constant.
@@
The form of Eq. 3 is still imperfect. Note that the gradient norm $\|\mathbf{g}\|$ decreases along with LR during the annealing process (shown in Fig. 4). Thus, the amount of parameter movement (approximately $\eta\cdot\|\mathbf{g}\|$ per step) in the LR annealing stage declines at an almost quadratic rate compared to stages before annealing. As the parameter movement become smaller, the change in loss also slows down accordingly. Therefore, the loss drop brought by the power law term (i.e., the first two terms in Eq. 3) should also diminish during LR annealing. This consideration leads to an improved equation:
@@
$S_{1}$ $=\sum\limits_{i=1}^{s}\eta_{i},$
where $S_{1}$ is the forward area, i.e., the area under the LR curve (as visualized in Fig. 1), which could be approximately interpreted as the total amount of parameter updates.